OpenStudy (anonymous):

Factor completely: x2 − 3x + 7 Prime (x + 7)(x − 1) (x − 7)(x − 1) (x + 7)(x + 1)

OpenStudy (anonymous):

@girlychix27

OpenStudy (anonymous):

Same as before. Trying multiplying out each of the answers using distribution.

OpenStudy (nincompoop):

quit cheating, @christianfilms11 try to learn how to do this one using the AC-Method

OpenStudy (anonymous):

I'm not cheating, I'm learning.

OpenStudy (anonymous):

Teach me please. @girlychix27

OpenStudy (anonymous):

(x + 7)(x − 1) = x2 -x + 7x - 7 which simplifies x2 + 7x - 7. so this is not a solution.

OpenStudy (anonymous):

(x − 7)(x − 1) = x2 -x -7x + 7 which simplifies to x2 -8x + 7. So this is also not a solution. Try to see if the last one is a solution. If not, then the answer is prime.

OpenStudy (anonymous):

How would I figure it out?

OpenStudy (anonymous):

That doesn't seam too hard I'll try and do the last one and can you tell me if I'm right?

OpenStudy (anonymous):

Sure thing.

OpenStudy (anonymous):

x+x1+7x+8 is this right?

OpenStudy (anonymous):

Almost. It would come out to x2 + x+ 7x + 8. Im not sure exactly where you got lost.

OpenStudy (anonymous):

Do you want me to tell you what I did?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

(x+7)(x+1) I added the first x with the second x and the 1 Then I added the 7 with the last x and the 1 Getting x+x1+7x+8

OpenStudy (anonymous):

Let me try to explain the foil method step by step. Just to clarify * means multiply (x+7)(x+1) 1) First you are going to multiply the "firsts." So in this case, x*x= x2. 2) Then you multiply the "outsides." So x*1 = x. 3) Then you multiply the "insides." so 7*x = 7x 4) Finally you do the "lasts." So 7*1 = 7 Now you add up all the products. So x2+x+7x+7 which when you simplify is x2+8x+7

OpenStudy (anonymous):

Oh I see, thanks for clearing that up for me.

OpenStudy (anonymous):

You're welcome.

OpenStudy (anonymous):

But what will the answer be?

OpenStudy (anonymous):

@girlychix27